# Binding and dimerization control phase separation in a compartment

### v1

@@ III.2 Dimerization facilitates phase separation
We next investigate how dimerization affects phase separation in the simple situation where the bulk is negligible ($\Delta\omega\rightarrow\infty$). Since we now allow for phase separation ($\chi\neq 0$), the compartment can be heterogenous and its state is described by the fields $\phi_{1}(\bm{r})$ and $\phi_{2}(\bm{r})$. However, the dimerization equilibrium, $\delta F/\delta\phi_{1}=\delta F/\delta\phi_{2}$, holds everywhere, allowing us to express these fields in terms of total fraction $\phi(\bm{r})$, analogously to Eq. (5); see App. A. Consequently, the total free energy is given by an integral over a free energy density that only depends on $\phi(\bm{r})$, which is analogous to standard binary phase separation [33, 34]. In particular, the compartment can separate into a dense and a dilute phase with identical dimerization equilibrium, $\phi_{2}/\phi_{1}^{2}=K_{\mathrm{d}}$, if the interaction parameter $\chi$ is sufficiently large. Fig. 3a shows that larger $\chi$ is required to induce phase separation for smaller $\omega_{2}$ (smaller $K_{\mathrm{d}}$). This effect can be understood by analyzing the homogeneous state $\phi(\bm{r})=\bar{\phi}$, which becomes unstable in the spinodal region, $\chi>\chi_{\mathrm{sp}}$, with
@@
see App. A. This expression shows that $\chi_{\mathrm{sp}}$ increases for smaller $K_{\mathrm{d}}=e^{2\omega_{2}+1}$. This effect is stronger for smaller $\bar{\phi}$, implying that smaller $\omega_{2}$ move the spinodal region toward large total fractions; see Fig. 3a. This behavior can also be understood from the limiting case of only monomers ($\omega_{2}\rightarrow-\infty$) and only dimers ($\omega_{2}\rightarrow\infty$): In terms of the Flory-Huggins theory [32] the only difference between these two cases is the contribution to translational entropy, which is halved for the dimers. Consequently, a smaller enthalpic gain by the unspecific interaction $\chi$ can offset the entropic loss of concentrating particles in a phase. The minimal value of $\chi$ necessary to obtain phase separation can be calculated by minimizing $\chi_{\mathrm{sp}}(\bar{\phi})$ given by Eq. (6), which corresponds to the critical point (stars in Fig. 3a). For monomers ($\omega_{2}\rightarrow-\infty$), we obtain $\chi_{\mathrm{sp}}^{\mathrm{min}}=2$, whereas $\chi_{\mathrm{sp}}^{\mathrm{min}}=(3+2\sqrt{2})/4\simeq 1.46$ for dimers ($\omega_{2}\rightarrow\infty$). Taken together, this analysis shows that the dimerization energy $\omega_{2}$ can control phase separation and droplet size; see Fig. 3b.
@@
Figure 3: Dimerization facilitates phase separation. (a) Phase diagrams as a function of volume fraction $\phi$ and interaction strength $\chi$ for the two indicated dimerization energies $\omega_{2}$. Phase-separated states are stable above the solid binodal lines, whereas the homogeneous state is stable below the dashed spinodal lines. The critical points are marked by stars. (b) Fraction of the volume $V^{\mathrm{II}}_{\mathrm{c}}$ of the dense phase to the compartment volume $V_{\mathrm{c}}$ as a function of $\omega_{2}$ for three different parameters indicated in the plot and marked in panel (a).
@@ III.3 Sequestration by the bulk suppresses phase separation in the compartment
We next study phase separation in the compartment coupled to the bulk without dimerization (only monomers, $\omega_{2}\rightarrow-\infty$). We focus on the case where particles prefer the compartment ($\Delta\omega>0$) since in the opposing case the smaller compartment would be negligible compared to the bulk. In this case, the average fraction $\bar{\phi}$ in the compartment is larger than the fraction $\bar{\psi}$ in the bulk, favoring phase separation in the compartment. The bulk would only phase separate if the total fraction $\Phi_{\mathrm{tot}}$ is very large, at which point the compartment is saturated. In this case, the compartment would only sequester a fixed amount of material, but not influence phase separation further. More generally, at most one of two coupled subsystems can exhibit phase separation in equilibrium [13, 14, 15, 30], and we thus focus on the case where the compartment is favored ($\Delta\omega>0$) and the bulk remains homogeneous with a fraction $\bar{\psi}$. In this case, the bulk can be approximated as a dilute mixture, so the chemical potentials in the compartment, $\mu_{\mathrm{c}}=\delta F/\delta\phi$, and the bulk, $\mu_{\mathrm{b}}=\partial F/\partial\bar{\psi}$, read
@@
$\mu_{\mathrm{c}}(\phi)$ $=\ln\left(\frac{\phi}{1-\phi}\right)-\Delta\omega+\chi(1-2\phi)\;,$ (7a)
$\mu_{\mathrm{b}}(\bar{\psi})$ $\approx\ln(\bar{\psi})+\chi\;,$ (7b)
@@
for $\bar{\psi}\ll 1$. Using material conservation given by Eq. (2) to eliminate $\bar{\psi}$ in Eq. (7b), and shifting both chemical potentials by $\ln\eta-\chi$, we find
@@
$\mu^{\mathrm{eff}}_{\mathrm{c}}(\phi)$ $=\ln\left(\frac{\phi}{1-\phi}\right)-\Delta\omega_{\mathrm{eff}}-2\phi\chi\;,$ (8a)
$\mu^{\mathrm{eff}}_{\mathrm{b}}(\bar{\phi})$ $\approx\ln(\Phi_{\mathrm{tot}}-\bar{\phi})\;,$ (8b)
@@
where $\Delta\omega_{\mathrm{eff}}=\Delta\omega-\ln\eta$ captures the enthalpy and entropy gained when a solute particle moves from the bulk to the compartment. It shows that the relative size $\eta$ of the subsystems plays a similar role to the compartment affinity $\Delta\omega$. In particular, in the limit of a small bulk ($\eta\rightarrow 0$), we find $\Delta\omega_{\mathrm{eff}}\rightarrow\infty$, so that all solute particles go to the compartment ($\bar{\phi}\rightarrow\Phi_{\mathrm{tot}}$). In this case, the bulk becomes negligible, and the behavior in the compartment is governed by ordinary phase separation in the canonical ensemble. In the contrasting case of a large bulk ($\eta\rightarrow\infty$), implying $\Delta\omega_{\mathrm{eff}}\rightarrow-\infty$, we also need to increase the total fraction $\Phi_{\mathrm{tot}}$ to have solutes in the compartment. This implies $\Phi_{\mathrm{tot}}\gg\bar{\phi}$, so that $\bar{\phi}$ is negligible in Eq. (8b), implying constant $\mu_{\mathrm{b}}$, which corresponds to a grand-canonical ensemble for the compartment. The interesting case of a bulk of finite size thus interpolates between these two traditional ensembles.
@@
Fig. 4a shows how the phase diagram depends on the effective affinity $\Delta\omega_{\mathrm{eff}}$. In the limit of large $\Delta\omega_{\mathrm{eff}}$ (blue lines), we indeed recover the canonical phase diagram of a binary system since the bulk is almost empty $\bar{\psi}\approx 0$, whereas the total fraction $\bar{\phi}$ in the compartment approaches the total fraction $\Phi_{\mathrm{tot}}$. Phase separation can thus take place in the binodal region, where $\Phi_{\mathrm{tot}}$ is between the fractions $\phi^{\mathrm{I}}$ and $\phi^{\mathrm{II}}$ of the dilute and dense phase in the compartment, $\Phi_{\mathrm{tot}}\in(\phi^{\mathrm{I}},\phi^{\mathrm{II}})$. For smaller $\Delta\omega_{\mathrm{eff}}$ the entire phase diagram shifts toward larger $\Phi_{\mathrm{tot}}$ (orange lines), essentially because a significant fraction of solute molecules is in the bulk. In particular, phase separation only occurs in the compartment when $\Phi_{\mathrm{tot}}\in(\Phi_{\mathrm{tot}}^{\mathrm{min}},\Phi_{\mathrm{tot}}^{\mathrm{max}})$; see Fig. 4a for an example at $\chi=3$. However, the coexisting fractions in the dilute and dense phase in the compartment are still given by $\phi^{\mathrm{I}}$ and $\phi^{\mathrm{II}}$ (see App. C), and thus need to be read off the phase diagram corresponding to the canonical ensemble (blue lines). In essence, the phase diagram as a function of $\Phi_{\mathrm{tot}}$ determines what total fractions lead to phase separation, whereas the coexisting volume fractions $\phi^{\mathrm{I/II}}$ need to be determined from the canonical phase diagram as a function of $\bar{\phi}$. We show in App. C that the difference between these fractions are
@@
because the bulk sequesters a constant fraction of material when the compartment phase separates; see Fig. 4b. In essence, phase coexistence in the compartment determines the fractions $\phi^{\mathrm{I/II}}$, with equal chemical potentials between the two phases and the bulk. Changing $\Phi_{\mathrm{tot}}$ then only changes the relative sizes of the two phases according to the lever rule, leaving all fractions unchanged. Eq. (9) shows that the fraction of material sequestered by the bulk decreases with larger effective affinity $\Delta\omega_{\mathrm{eff}}$ and larger interaction strength $\chi$, so that the phase diagram approaches the canonical limit for large $\Delta\omega_{\mathrm{eff}}+\chi$.
@@
Bulk sequestration changes the phase diagram qualitatively. For instance, we find a re-entrant phase transition when changing the interaction parameter $\chi$. Fig. 4a shows an example for $\Delta\omega_{\mathrm{eff}}=-2$ (orange lines) where phase separation is only possible in a narrow regime of $\chi$ when $\Phi_{\mathrm{tot}}$ is equal to the marked $\Phi_{\mathrm{tot}}^{\mathrm{max}}$. This is essentially due to an increasing fraction $\bar{\phi}$ in the compartment when $\chi$ increases, caused by the lower chemical potential; see Eq. (8a). Similarly, $\bar{\phi}$ increases monotonically (and $\bar{\psi}$ decreases accordingly) for larger affinity $\Delta\omega$ or smaller size ratio $\eta$. In contrast, the total fraction $\Phi_{\mathrm{tot}}$ affects the system differently; see Fig. 4b. While $\bar{\phi}$ and $\bar{\psi}$ increase with $\Phi_{\mathrm{tot}}$ if the compartment is homogeneous (outside the binodal), only the size of the dense phase increases, but the fractions stay constant, when the compartment phase separates, which is akin to noise buffering described in simple phase separation [35]. Taken together, we show that the affinity $\Delta\omega$, the size ratio $\eta$, and the interaction strength $\chi$ can be used to control phase separation in the compartment together with the total amount $\Phi_{\mathrm{tot}}$.
@@
We next discuss the stability of homogeneous states to distinguish nucleation and growth regimes from spinodal decomposition. Fig. 4c shows that the spinodal (dashed orange line) also moves toward larger total fractions $\Phi_{\mathrm{tot}}$ for smaller $\Delta\omega_{\mathrm{eff}}$, similarly to the binodals discussed above. In this case, the entire phase diagram is thus qualitatively similar to the phase diagram in the canonical limit ($\Delta\omega_{\mathrm{eff}}\rightarrow\infty$). In contrast, phase diagrams for smaller $\Delta\omega_{\mathrm{eff}}$, where the compartment is less favored, can be qualitatively different; see Fig. 4d. In this case, the spinodal decomposition region is reduced dramatically, and the two nucleation-and-growth regimes overlap. In this overlap region (dark green), there are three stable states: the phase separated state has the lowest free energy, but there are also two locally stable homogeneous states. This is because the binding equilibrium, $\mu_{\mathrm{c}}(\bar{\phi})=\mu_{\mathrm{b}}(\bar{\phi})$ using Eqs. (8), leads to a non-linear equation with three solutions, of which two are stable; see App. C. We thus showed that a medium-sized bulk can affect phase separation qualitatively, leading to the effective binding affinity $\Delta\omega_{\mathrm{eff}}=\Delta\omega-\ln\eta$ as a central control parameter.
@@
Figure 4: Sequestration by the bulk suppresses phase separation in the compartment. (a) Phase diagrams as a function of the total fraction $\Phi_{\mathrm{tot}}$ and interaction strength $\chi$ for two effective affinities $\Delta\omega_{\mathrm{eff}}=\Delta\omega-\ln\eta$. The compartment can phase separate within the colored regions. The limit of strong binding (blue) recovers the canonical phase diagram, determining the coexisting volume fractions $\phi^{\mathrm{I}}$ and $\phi^{\mathrm{II}}$ (plot shows example for $\chi=3$). Weaker binding (green and orange) moves the region to larger $\Phi_{\mathrm{tot}}$ and the intersections with the line $\chi=3$ now give the region $\Phi_{\mathrm{tot}}\in(\Phi_{\mathrm{tot}}^{\mathrm{min}},\Phi_{\mathrm{tot}}^{\mathrm{max}})$ where phase separation takes place, whereas the coexisting volume fractions still correspond to $\phi^{\mathrm{I/II}}$. (b) Average volume fractions $\bar{\phi}$ in the compartment and normalized fraction $\eta\bar{\psi}$ in the bulk as a function of $\Phi_{\mathrm{tot}}$ for $\Delta\omega_{\mathrm{eff}}=-2$. (c, d) Phase diagrams including spinodals (dashed lines) for two values of $\Delta\omega_{\mathrm{eff}}$, separating spinodal decomposition (SD) from the nucleation-and-growth (NG) region. White regions denote a globally stable homogeneous configuration (HC), whereas HCs are only linearly stable in the darker region between the dashed and solid lines, leading to multistability in the central region of panel (d). The bulk exhibits phase separation in the gray region for $\eta=100$.

### v2

@@ III.2 Dimerization facilitates phase separation
We next investigate how dimerization affects phase separation in the simple situation where the bulk is negligible ($\Delta\omega\rightarrow\infty$), so we only focus on the compartment. Since we now allow for phase separation ($\chi\neq 0$), the compartment can be heterogenous and its state is described by the fields $\phi_{1}(\bm{r})$ and $\phi_{2}(\bm{r})$. However, the dimerization equilibrium, $\delta F/\delta\phi_{1}=\delta F/\delta\phi_{2}$, holds everywhere, allowing us to express these fields in terms of solute fraction $\phi(\bm{r})$, analogously to Eq. (5); see Appendix B. Consequently, the total free energy is given by an integral over a free energy density that only depends on $\phi(\bm{r})$, which is analogous to standard binary phase separation [35, 36]. In particular, the compartment can separate into a dense and a dilute phase with identical dimerization equilibrium, $\phi_{2}/\phi_{1}^{2}=K_{\mathrm{d}}$, if the interaction parameter $\chi$ is sufficiently large. Fig. 3a shows that larger $\chi$ is required to induce phase separation for smaller $\omega_{2}$ (smaller $K_{\mathrm{d}}$). This effect can be understood by analyzing the homogeneous state $\phi(\bm{r})=\bar{\phi}$, which becomes unstable in the spinodal region, $\chi>\chi_{\mathrm{sp}}$, with
@@
see Appendix B. This expression shows that $\chi_{\mathrm{sp}}$ increases for smaller $K_{\mathrm{d}}=e^{2\omega_{2}+1}$. This effect is stronger for smaller $\bar{\phi}$, implying that smaller $\omega_{2}$ move the spinodal region toward large solute fractions; see Fig. 3a. This behavior can also be understood from the limiting case of only monomers ($\omega_{2}\rightarrow-\infty$) and only dimers ($\omega_{2}\rightarrow\infty$): In terms of the Flory-Huggins theory [34] the only difference between these two cases is the contribution to translational entropy, which is halved for the dimers. Consequently, a smaller enthalpic gain by the unspecific interaction $\chi$ can offset the entropic loss of concentrating particles in a phase. The minimal value of $\chi$ necessary to obtain phase separation can be calculated by minimizing $\chi_{\mathrm{sp}}(\bar{\phi})$ given by Eq. (6), which corresponds to the critical point (stars in Fig. 3a). For monomers ($\omega_{2}\rightarrow-\infty$), we obtain $\chi_{\mathrm{sp}}^{\mathrm{min}}=2$, whereas $\chi_{\mathrm{sp}}^{\mathrm{min}}=(3+2\sqrt{2})/4\simeq 1.46$ for dimers ($\omega_{2}\rightarrow\infty$). Taken together, this analysis shows that the dimerization energy $\omega_{2}$ can control phase separation and droplet size; see Fig. 3b.
@@
Figure 3: Dimerization facilitates phase separation. (a) Phase diagrams as a function of solute fraction $\bar{\phi}$ and interaction strength $\chi$ for the two indicated dimerization energies $\omega_{2}$. Phase-separated states are stable above the solid binodal lines, whereas the homogeneous state is stable below the dashed spinodal lines. The critical points are marked by stars. (b) Fraction of volume $V^{\mathrm{II}}_{\mathrm{c}}$ of the dense phase relative to the compartment volume $V_{\mathrm{c}}$ as a function of $\omega_{2}$ for various $\chi$ and $\bar{\phi}$ indicated in the plot and marked in panel a. Vertical lines correspond to the values of $\omega_{2}$ shown in panel a.
@@ III.3 Sequestration by the bulk suppresses phase separation in the compartment
We next study phase separation in the compartment coupled to the bulk without dimerization (only monomers, $\omega_{2}\rightarrow-\infty$). We focus on the case where particles prefer the compartment ($\Delta\omega>0$) since in the opposing case the smaller compartment would be negligible compared to the bulk. In this case, the average fraction $\bar{\phi}$ in the compartment is larger than the fraction $\bar{\psi}$ in the bulk, favoring phase separation in the compartment. The bulk would only phase separate if the total fraction $\Phi_{\mathrm{tot}}$ is very large, at which point the compartment is saturated. In this case, the compartment would only sequester a fixed amount of material, but not influence phase separation further. More generally, at most one of two coupled subsystems can exhibit phase separation in equilibrium [14, 15, 16, 31]. This results from Gibbs phase rule [37], which allows at most two phases per subsystem if they are not in contact; enabling material exchange (but keeping the volumes fixed) adds one constraint, so the total phase count is three and not four.
Since only one subsystem can exhibit phase separation, we focus on the case where the compartment is favored ($\Delta\omega>0$) and the bulk remains homogeneous with a fraction $\bar{\psi}$. In this case, the bulk can be approximated as a dilute mixture, so the chemical potentials in the compartment, $\mu_{\mathrm{c}}=\nu_{\mathrm{c}}\partial f_{\mathrm{c}}/\partial\phi$, and the bulk, $\mu_{\mathrm{b}}=\nu_{\mathrm{b}}\partial f_{\mathrm{b}}/\partial\bar{\psi}$, read
@@
$\mu_{\mathrm{c}}(\phi)$ $=k_{\mathrm{B}}T\Bigl{[}\ln\left(\frac{\phi}{1-\phi}\right)-\Delta\omega+\chi(1-2\phi)\Bigr{]}\;,$ (7a)
$\mu_{\mathrm{b}}(\bar{\psi})$ $\approx k_{\mathrm{B}}T\Bigl{[}\ln(\bar{\psi})+\chi\Bigr{]}\;,$ (7b)
@@
for $\bar{\psi}\ll 1$. Since the assumption that the bulk is dilute could be violated in some cases, we will check it a posteriori. Using material conservation given by Eq. (2) to eliminate $\bar{\psi}$ in Eq. (7b), and shifting both chemical potentials by $\ln\eta-\chi$, we find
@@
$\mu^{\mathrm{eff}}_{\mathrm{c}}(\phi)$ $=k_{\mathrm{B}}T\Bigl{[}\ln\left(\frac{\phi}{1-\phi}\right)-\Delta\omega_{\mathrm{eff}}-2\phi\chi\Bigr{]}\;,$ (8a)
$\mu^{\mathrm{eff}}_{\mathrm{b}}(\bar{\phi})$ $\approx k_{\mathrm{B}}T\Bigl{[}\ln(\Phi_{\mathrm{tot}}-\bar{\phi})\Bigr{]}\;,$ (8b)
@@
where $\Delta\omega_{\mathrm{eff}}=\Delta\omega-\ln\eta$ captures the enthalpy and entropy gained when a solute particle moves from the bulk to the compartment. It shows that the relative size $\eta$ of the subsystems plays a similar role to the compartment affinity $\Delta\omega$. In particular, in the limit of a small bulk ($\eta\rightarrow 0$), we find $\Delta\omega_{\mathrm{eff}}\rightarrow\infty$, so that all solute particles go to the compartment ($\bar{\phi}\rightarrow\Phi_{\mathrm{tot}}$). In this case, the bulk becomes negligible, the total mass in the compartment is conserved, and the behavior in the compartment is governed by ordinary phase separation in the canonical ensemble. In the contrasting case of a large bulk ($\eta\rightarrow\infty$), implying $\Delta\omega_{\mathrm{eff}}\rightarrow-\infty$, we also need to increase the total fraction $\Phi_{\mathrm{tot}}$ to have solutes in the compartment. This implies $\Phi_{\mathrm{tot}}\gg\bar{\phi}$, so that $\bar{\phi}$ is negligible in Eq. (8b), implying constant $\mu_{\mathrm{b}}$, which corresponds to a grand-canonical ensemble for the compartment. The interesting case of a bulk of finite size thus interpolates between these two traditional ensembles.
@@
Fig. 4a shows how the phase diagram depends on the effective affinity $\Delta\omega_{\mathrm{eff}}$. In the limit of large $\Delta\omega_{\mathrm{eff}}$ (blue lines), we indeed recover the canonical phase diagram of a binary system since the bulk is almost empty $\bar{\psi}\approx 0$, whereas the solute fraction $\bar{\phi}$ in the compartment approaches the total fraction $\Phi_{\mathrm{tot}}$. Phase separation can thus take place in the binodal region, where $\Phi_{\mathrm{tot}}$ is between the fractions $\phi^{\mathrm{I}}$ and $\phi^{\mathrm{II}}$ of the dilute and dense phase in the compartment, $\Phi_{\mathrm{tot}}\in(\phi^{\mathrm{I}},\phi^{\mathrm{II}})$. For smaller $\Delta\omega_{\mathrm{eff}}$ the entire phase diagram shifts toward larger $\Phi_{\mathrm{tot}}$ (orange lines), essentially because a significant fraction of solute molecules are in the bulk. In particular, phase separation only occurs in the compartment when $\Phi_{\mathrm{tot}}\in(\Phi_{\mathrm{tot}}^{\mathrm{min}},\Phi_{\mathrm{tot}}^{\mathrm{max}})$; see Fig. 4a for an example at $\chi=3$. However, the coexisting fractions in the dilute and dense phase in the compartment are still given by $\phi^{\mathrm{I}}$ and $\phi^{\mathrm{II}}$ (see Appendix D), and thus need to be read off the phase diagram corresponding to the canonical ensemble (blue lines). In essence, the phase diagram as a function of $\Phi_{\mathrm{tot}}$ determines what total fractions lead to phase separation, whereas the coexisting volume fractions $\phi^{\mathrm{I/II}}$ need to be determined from the canonical phase diagram as a function of $\bar{\phi}$. We show in Appendix D that the difference between these fractions are
@@
because the bulk sequesters a constant fraction of material when the compartment phase separates; see Fig. 4b. The fraction of material sequestered by the bulk thus decreases with larger effective affinity $\Delta\omega_{\mathrm{eff}}$ and larger interaction strength $\chi$, so that the phase diagram approaches the canonical limit for large $\Delta\omega_{\mathrm{eff}}+\chi$. In essence, $\Delta\omega$ and $\chi$ determine how much material ends up in the compartment, where phase coexistence then determines the fractions $\phi^{\mathrm{I/II}}$. Changing $\Phi_{\mathrm{tot}}$ then only changes the relative sizes of the two phases according to the lever rule, leaving all fractions unchanged. The fraction of volume of the dense phase reads
$\frac{V_{\mathrm{c}}^{\mathrm{II}}}{V_{\mathrm{c}}}=\frac{1}{\phi^{\mathrm{II}}-\phi^{\mathrm{I}}}\Bigl{(}\Phi_{\text{tot}}-\phi^{\mathrm{I}}-e^{-\Delta\omega_{\mathrm{eff}}-\chi}\Bigr{)}\;,$ (10)
showing that the dense phase in the compartment increases when there is more material (larger $\Phi_{\mathrm{tot}}$) and larger effective affinity $\Delta\omega_{\mathrm{eff}}$.
@@
Bulk sequestration changes the phase diagram qualitatively. For instance, we find a re-entrant phase transition when changing the interaction parameter $\chi$. Fig. 4a shows an example for $\Delta\omega_{\mathrm{eff}}=-2$ (orange lines) where phase separation is only possible in a narrow regime of $\chi$ when $\Phi_{\mathrm{tot}}$ is equal to the marked $\Phi_{\mathrm{tot}}^{\mathrm{max}}$. This is essentially due to an increasing fraction $\bar{\phi}$ in the compartment when $\chi$ increases, caused by the lower chemical potential; see Eq. (8a). Similarly, $\bar{\phi}$ increases monotonically (and $\bar{\psi}$ decreases accordingly) for larger affinity $\Delta\omega$ or smaller size ratio $\eta$. In contrast, the total fraction $\Phi_{\mathrm{tot}}$ affects the system differently; see Fig. 4b. While $\bar{\phi}$ and $\bar{\psi}$ increase with $\Phi_{\mathrm{tot}}$ if the compartment is homogeneous (outside the binodal), only the size of the dense phase increases, but the fractions stay constant, when the compartment phase separates, which is akin to noise buffering described in simple phase separation [38]. Taken together, we show that the affinity $\Delta\omega$, the size ratio $\eta$, and the interaction strength $\chi$ can be used to control phase separation in the compartment together with the total fraction $\Phi_{\mathrm{tot}}$.
@@
We next discuss the stability of homogeneous states to distinguish nucleation and growth regimes from spinodal decomposition. Fig. 4c shows that the spinodal (dashed orange line) also moves toward larger total fractions $\Phi_{\mathrm{tot}}$ for smaller $\Delta\omega_{\mathrm{eff}}$, similarly to the binodals discussed above. In this case, the entire phase diagram is thus qualitatively similar to the phase diagram in the canonical limit ($\Delta\omega_{\mathrm{eff}}\rightarrow\infty$). In contrast, phase diagrams for smaller $\Delta\omega_{\mathrm{eff}}$, where the compartment is less favored, can be qualitatively different; see Fig. 4d. In this case, the spinodal decomposition region is reduced dramatically, and the two nucleation-and-growth regimes overlap. In this overlap region (dark green), there are three stable states: the phase separated state has the lowest free energy, but there are also two locally stable homogeneous states. This is because the binding equilibrium, $\mu_{\mathrm{c}}(\bar{\phi})=\mu_{\mathrm{b}}(\bar{\phi})$ using Eqs. (8), leads to a non-linear equation with three solutions, of which two are stable; see Appendix D. Note that the bulk can also exhibit phase separation when the compartment is densely occupied (at large $\Phi_{\mathrm{tot}}$ and $\chi$; gray region in Fig. 4), but we do not analyze this case further since it is roughly equivalent to phase separation in the bulk with part of the material sequestered by the compartment. In the more interesting case where the compartment exhibits phase separation, we showed that a medium-sized bulk can affect phase separation qualitatively, leading to the effective binding affinity $\Delta\omega_{\mathrm{eff}}=\Delta\omega-\ln\eta$ as a central control parameter.
@@
Figure 4: Sequestration by the bulk suppresses phase separation in the compartment. (a) Phase diagrams as a function of the total fraction $\Phi_{\mathrm{tot}}$ and interaction strength $\chi$ for two effective affinities $\Delta\omega_{\mathrm{eff}}=\Delta\omega-\ln\eta$. The compartment can phase separate within the colored regions. The limit of strong binding (blue) recovers the canonical phase diagram, determining the coexisting volume fractions $\phi^{\mathrm{I}}$ and $\phi^{\mathrm{II}}$ (plot shows example for $\chi=3$). Weaker binding (green and orange) moves the region to larger $\Phi_{\mathrm{tot}}$ and the intersections with the line $\chi=3$ now give the region $\Phi_{\mathrm{tot}}\in(\Phi_{\mathrm{tot}}^{\mathrm{min}},\Phi_{\mathrm{tot}}^{\mathrm{max}})$ where phase separation takes place, whereas the coexisting volume fractions still correspond to $\phi^{\mathrm{I/II}}$. (b) Average volume fractions $\bar{\phi}$ in the compartment and normalized fraction $\eta\bar{\psi}$ in the bulk as a function of $\Phi_{\mathrm{tot}}$ for $\Delta\omega_{\mathrm{eff}}=-2$ and $\chi=3$. (c, d) Phase diagrams including spinodals (dashed lines) for two values of $\Delta\omega_{\mathrm{eff}}$, separating spinodal decomposition (SD) from the nucleation-and-growth (NG) region. White regions denote a globally stable homogeneous configuration (HC), whereas HCs are only linearly stable in the darker region between the dashed and solid lines, leading to multistability in the central region of panel (d). The bulk exhibits phase separation in the gray region for $\eta=100$.